OpenStudy (anonymous):

The average of 11 numbers is 21 . The average of seven of those numbers is 25. What is the average of the remaining four numbers?

OpenStudy (anonymous):

28

OpenStudy (anonymous):

my medal plz

OpenStudy (anonymous):

how did u get that

OpenStudy (anonymous):

look at the difference from 11-7 u go down by four digits which u then add to the next average,then do the same for 7-4 go down by 3 digits then add to the last average

OpenStudy (anonymous):

can you give me an equation example

OpenStudy (phi):

The trick is to know that to find the average of N numbers, you add them up and divide by N. Say you have 11 numbers called a,b,c,d,e,f,g,h,i,j,k; and their average is 21. This means \[\frac{a+b+c+d+e+f+g+h+i+j+k}{11}=21\] and, by multiplying both sides of the equation by 11: \[a+b+c+d+e+f+g+h+i+j+k=21*11= 231\] You also know seven of the numbers, say a through g average 25. This means \[a+b+c+d+e+f+g=25*7= 175\] if you subtract the bottom equation from the top equation, you get \[h+i+j+k=56\] So you know how much the 4 numbers add up to. Divide by 4 to get their average. \[\frac{h+i+j+k}{4}=\frac{56}{4}= 14\]